# GaussianAnything: Interactive Point Cloud Latent Diffusion for 3D Generation

### v1

@@ Abstract
While 3D content generation has advanced significantly, existing methods still face challenges with input formats, latent space design, and output representations. This paper introduces a novel 3D generation framework that addresses these challenges, offering scalable, high-quality 3D generation with an interactive Point Cloud-structured Latent space. Our framework employs a Variational Autoencoder (VAE) with multi-view posed RGB-D(epth)-N(ormal) renderings as input, using a unique latent space design that preserves 3D shape information, and incorporates a cascaded latent diffusion model for improved shape-texture disentanglement. The proposed method, GaussianAnything, supports multi-modal conditional 3D generation, allowing for point cloud, caption, and single/multi-view image inputs. Notably, the newly proposed latent space naturally enables geometry-texture disentanglement, thus allowing 3D-aware editing. Experimental results demonstrate the effectiveness of our approach on multiple datasets, outperforming existing methods in both text- and image-conditioned 3D generation.
@@ 1 Introduction
3D content generation holds great potential for transforming the virtual reality, film, and gaming industries. Current approaches typically follow one of two paths: either a 2D-lifting method or the design of native 3D diffusion models. While the 2D-lifting approach (Shi et al., 2023b; Liu et al., 2023b) benefits from leveraging 2D diffusion model priors, it is often hindered by expensive optimization, the Janus problem, and inconsistencies between views. In contrast, native 3D diffusion models (Jun & Nichol, 2023; Lan et al., 2024; Zhang et al., 2024) are trained from scratch for 3D generation, offering improved generality, efficiency, and control.
@@
Despite the progress in native 3D diffusion models, several design challenges still persist: (1) Input format to the 3D VAE. Most methods (Zhang et al., 2024; Li et al., 2024) directly adopt point cloud as input. However, it fails to encode the high-frequency details from textures. Besides, this limits the available training dataset to artist-created 3D assets, which are challenging to collect on a large scale. LN3Diff (Lan et al., 2024) adopt multi-view images as input. Though straightforward, it lacks direct 3D information input and cannot comprehensively encode the given object. (2) 3D latent space structure. Since 3D objects are diverse in geometry, color, and size, most 3D VAE models adopt the permutation-invariant set latent (Zhang et al., 2023a; Sajjadi et al., 2022; Zhang et al., 2024) to encode incoming 3D objects. Though flexible, this design lacks the image-latent correspondence as in Stable Diffusion VAE (Rombach et al., 2022), where the VAE latent code can directly serve as the proxy for editing input image (Mou et al., 2023b; a). Other methods adopt latent tri-plane (Wu et al., 2024; Lan et al., 2024) as the 3D latent representation. However, the latent tri-plane is still unsuitable for interactive 3D editing as changes in one plane may not map to the exact part of the objects that need editing. (3) Choice of 3D output representations. Existing solutions either output texture-less SDF (Wu et al., 2024; Zhang et al., 2024), which requires additional shading model for post-processing; or volumetric tri-plane (Lan et al., 2024), which struggles with high-resolution rendering due to extensive memory required by volumetric rendering (Mildenhall et al., 2020).
@@
In this study, we propose a novel 3D generation framework that resolves the problems above and enables scalable, high-quality 3D generation with an interactive Point Cloud-structured Latent space. The resulting method, dubbed GaussianAnything, supports multi-modal conditional 3D generation, including point cloud, caption, and image. Specifically, we propose a 3D VAE that adopts multi-view posed RGB-D(epth)-N(ormal) renderings as the input, which are easy to render and contain comprehensive 3D attributes corresponding to the input 3D object. The information of each input view is channel-wise concatenated and efficiently encoded with the scene representation transformer (Sajjadi et al., 2022), yielding a set latent that compactly encodes the given 3D input. Instead of directly applying it for diffusion learning (Zhang et al., 2024; Li et al., 2024), our novel design concretizes the unordered tokens into the shape of the 3D input. Specifically, this is achieved by cross-attending (Huang et al., 2024b) the set latent via a sparse point cloud sampled from the input 3D shape, as visualized in Fig. 2. The resulting point-cloud structured latent space significantly facilitate shape-texture disentanglement and 3D editing. Afterward, a DiT-based 3D decoder (Peebles & Xie, 2023; Lan et al., 2024) gradually decodes and upsamples the latent point cloud into a set of dense surfel Gaussians (Huang et al., 2024a), which are rasterized to high-resolution renderings to supervise 3D VAE training.
@@ 2 Related Work
3D Generation via 2D Diffusion Models. The success of 2D diffusion models (Song et al., 2021; Ho et al., 2020) has inspired their application to 3D generation. Score distillation sampling (Poole et al., 2022; Wang et al., 2023) distills 3D from a 2D diffusion model, but faces challenges like expensive optimization, mode collapse, and the Janus problem. More recent methods propose learning the 3D via a two-stage pipeline: multi-view images generation (Shi et al., 2023b; Long et al., 2024; Shi et al., 2023a) and feed-forward 3D reconstruction (Hong et al., 2024b; Xu et al., 2024; Tang et al., 2024). Though promising results have been achieved, their performance is bounded by the multi-view generation results, which usually violate view consistency (Liu et al., 2023b) and fails to scale up to higher resolution (Shi et al., 2023a). Moreover, this two-stage pipeline limits the 3D editing capability due to the lack of a 3D-aware latent space.
@@
Native 3D Diffusion Models. Native 3D diffusion models (Zhang et al., 2023a; Zeng et al., 2022; Zhang et al., 2024; Lan et al., 2024; Li et al., 2024) are recently proposed to achieve high-quality, efficient and scalable 3D generation. A native 3D diffusion pipeline involves a two-stage training process: encoding 3D objects into the VAE latent space (Kingma & Welling, 2013; Kosiorek et al., 2021), and latent diffusion model on the corresponding latent codes. Though straightforward, existing methods differ in VAE input formats, latent space structure and output 3D representations. While most methods adopt point alone as the VAE input (Zhang et al., 2023a; 2024; Li et al., 2024), our proposed method encodes a hybrid 3D information through convolutional encoder. Moreover, comparing to the latent set (Zhang et al., 2023a; Sajjadi et al., 2022) representation, our proposed method adopts a point cloud-structured latent space, which can be directly used for interactive 3D editing. Besides, rather than producing textureless SDF, our method directly decodes the 3D latent codes into high-quality surfel Gaussians (Huang et al., 2024a), which can be directly used for efficient rendering.
@@
Point-based Shape Representation and Rendering. The proliferation of 3D scanners and RGB-D cameras makes the capture and processing of 3D point clouds commonplace (Gross & Pfister, 2011). In the era of deep learning, learning-based methods are emerging for point set processing (Qi et al., 2016; Zhao et al., 2021), up-sampling (Yu et al., 2018), shape representation (Genova et al., 2020; Lan et al., 2023b), and rendering (Pfister et al., 2000; Yifan et al., 2019; Lassner & Zollhöfer, 2021; Xu et al., 2022; Kerbl et al., 2023). Moreover, given its affinity for modern network architectures (Huang et al., 2024b; Zhao et al., 2021), more explicit nature against other 3D representations (Chan et al., 2022; Mildenhall et al., 2020; Müller et al., 2022), efficient rendering (Kerbl et al., 2023), and even high-quality surface modeling (Huang et al., 2024a), point-based 3D representations are rapidly developing towards the canonical 3D representation for learning 3D shapes. Thus, we choose (featured) point cloud as the representation for the 3D VAE latent space, and 2D Gaussians (Huang et al., 2024a) as the output 3D representations.

### v2

# GaussianAnything: Interactive Point Cloud Flow Matching for 3D Object Generation

@@ Abstract
While 3D content generation has advanced significantly, existing methods still face challenges with input formats, latent space design, and output representations. This paper introduces a novel 3D generation framework that addresses these challenges, offering scalable, high-quality 3D generation with an interactive Point Cloud-structured Latent space. Our framework employs a Variational Autoencoder (VAE) with multi-view posed RGB-D(epth)-N(ormal) renderings as input, using a unique latent space design that preserves 3D shape information, and incorporates a cascaded latent flow-based model for improved shape-texture disentanglement. The proposed method, GaussianAnything, supports multi-modal conditional 3D generation, allowing for point cloud, caption, and single image inputs. Notably, the newly proposed latent space naturally enables geometry-texture disentanglement, thus allowing 3D-aware editing. Experimental results demonstrate the effectiveness of our approach on multiple datasets, outperforming existing native 3D methods in both text- and image-conditioned 3D generation.
@@ 1 Introduction
3D content generation holds great potential for transforming the virtual reality, film, and gaming industries. Current approaches typically follow one of two paths: either a 2D-lifting method or the design of native 3D diffusion models. While the 2D-lifting approach (Shi et al., 2023b; Liu et al., 2023b) benefits from leveraging 2D diffusion model priors, it is often hindered by expensive optimization, the Janus problem, and inconsistencies between views. In contrast, native 3D diffusion models (Jun & Nichol, 2023; Lan et al., 2024a; Zhang et al., 2024b) are trained from scratch for 3D generation, offering improved generality, efficiency, and control.
@@
Despite the progress in native 3D diffusion models, several design challenges still persist: (1) Input format to the 3D VAE. Most methods (Zhang et al., 2024b; Li et al., 2024) directly adopt point cloud as input. However, it fails to encode the high-frequency details from textures. Besides, this limits the available training dataset to artist-created 3D assets, which are challenging to collect on a large scale. LN3Diff (Lan et al., 2024a) adopt multi-view images as input. Though straightforward, it lacks direct 3D information input and cannot comprehensively encode the given object. (2) 3D latent space structure. Since 3D objects are diverse in geometry, color, and size, most 3D VAE models adopt the permutation-invariant set latent (Zhang et al., 2023; Sajjadi et al., 2022; Zhang et al., 2024b) to encode incoming 3D objects. Though flexible, this design lacks the image-latent correspondence as in Stable Diffusion VAE (Rombach et al., 2022), where the VAE latent code can directly serve as the proxy for editing input image (Mou et al., 2023b; a). Other methods adopt latent tri-plane (Wu et al., 2024; Lan et al., 2024a) as the 3D latent representation. However, the latent tri-plane is still unsuitable for interactive 3D editing as changes in one plane may not map to the exact part of the objects that need editing. (3) Choice of 3D output representations. Existing solutions either output texture-less SDF (Wu et al., 2024; Zhang et al., 2024b), which requires additional shading model for post-processing; or volumetric tri-plane (Lan et al., 2024a), which struggles with high-resolution rendering due to extensive memory required by volumetric rendering (Mildenhall et al., 2020).
@@
In this study, we propose a novel 3D generation framework that resolves the problems above and enables scalable, high-quality 3D generation with an interactive Point Cloud-structured Latent space. The resulting method, dubbed GaussianAnything, supports multi-modal conditional 3D generation, including point cloud, caption, and image. Specifically, we propose a 3D VAE that adopts multi-view posed RGB-D(epth)-N(ormal) renderings as the input, which are easy to render and contain comprehensive 3D attributes corresponding to the input 3D object. The information of each input view is channel-wise concatenated and efficiently encoded with the scene representation transformer (Sajjadi et al., 2022), yielding a set latent (Zhang et al., 2023) that compactly encodes the given 3D input. Instead of directly applying it for diffusion learning (Zhang et al., 2024b; Li et al., 2024), our novel design concretizes the unordered tokens into the shape of the 3D input. Specifically, this is achieved by cross-attending (Huang et al., 2024b) the set latent via a sparse point cloud sampled from the input 3D shape, as visualized in Fig. 1. The resulting point-cloud structured latent space significantly facilitate shape-texture disentanglement and 3D editing. Afterward, a DiT-based 3D decoder (Peebles & Xie, 2023; Lan et al., 2024a) gradually decodes and upsamples the latent point cloud into a set of dense surfel Gaussians (Huang et al., 2024a), which are rasterized to high-resolution renderings to supervise 3D VAE training.
@@ 2 Related Work
3D Generation via 2D Diffusion Models. The success of 2D diffusion models (Song et al., 2021; Ho et al., 2020) has inspired their application to 3D generation. Score distillation sampling (Poole et al., 2022; Wang et al., 2023) distills 3D from a 2D diffusion model, but faces challenges like expensive optimization, mode collapse, and the Janus problem. More recent methods propose learning the 3D via a two-stage pipeline: multi-view images generation (Shi et al., 2023b; Long et al., 2024; Shi et al., 2023a) and feed-forward 3D reconstruction and editing (Hong et al., 2024b; Xu et al., 2024a; Tang et al., 2024; Xu et al., 2024b; Chen et al., 2024b). However, their performance is bounded by the multi-view generation results, which usually violate view consistency (Liu et al., 2023b) and fail to scale up to higher resolution (Shi et al., 2023a). Moreover, this two-stage pipeline limits the 3D editing capability due to the lack of a 3D-aware latent space.
@@
Native 3D Diffusion Models. Native 3D diffusion models (Zhang et al., 2023; Zeng et al., 2022; Zhang et al., 2024b; Lan et al., 2024a; Li et al., 2024) are recently proposed to achieve high-quality, efficient and scalable 3D generation. A native 3D diffusion pipeline involves a two-stage training process: encoding 3D objects into the VAE latent space (Kingma & Welling, 2013; Kosiorek et al., 2021), and latent diffusion model on the corresponding latent codes. Though straightforward, existing methods differ in VAE input formats, latent space structure and output 3D representations. While most methods adopt point alone as the VAE input (Zhang et al., 2023; 2024b; Li et al., 2024), our proposed method encodes a hybrid 3D information through convolutional encoder. Moreover, comparing to the latent set (Zhang et al., 2023; Sajjadi et al., 2022) representation, our proposed method adopts a point cloud-structured latent space, which can be directly used for interactive 3D editing. Besides, rather than producing textureless SDF, our method directly decodes the 3D latent codes into high-quality surfel Gaussians (Huang et al., 2024a), which can be directly used for efficient rendering.
@@
Point-based Shape Representation and Rendering. The proliferation of 3D scanners and RGB-D cameras makes the capture and processing of 3D point clouds commonplace (Gross & Pfister, 2011). In the era of deep learning, learning-based methods are emerging for point set processing (Qi et al., 2016; Zhao et al., 2021), up-sampling (Yu et al., 2018), shape representation (Genova et al., 2020; Lan et al., 2024b), and rendering (Pfister et al., 2000; Yifan et al., 2019; Lassner & Zollhöfer, 2021; Xu et al., 2022; Kerbl et al., 2023). Moreover, given its affinity for modern network architectures (Huang et al., 2024b; Zhao et al., 2021), more explicit nature against other 3D representations (Chan et al., 2022; Mildenhall et al., 2020; Müller et al., 2022), efficient rendering (Kerbl et al., 2023), and even high-quality surface modeling (Huang et al., 2024a), point-based 3D representations are rapidly developing towards the canonical 3D representation for learning 3D shapes. Thus, we choose (featured) point cloud as the representation for the 3D VAE latent space, and 2D Gaussians (Huang et al., 2024a) as the output 3D representations.